A steady current of 10 A maintained for 1 hr deposits 12.2 g of zinc at the

cathode. What is the equivalent mass of zinc?


1
Expert's answer
2022-03-28T14:24:38-0400

We know that

"m=Zit"

"Z=\frac{molar mass}{nf\times F}"

"Z=\frac{65.4}{2\times96500}=3.4\times10^{-4}"


"m=3.4\times10^{-4}\times10\times1\times60\times60=12.24g"